# Low-Latency Video Conferencing System for Geo-Distributed Data Centers

### v1

@@ V Watermark-based Jitter Manager
The jitter manager sits on packet receivers’ side to ensure ordered emit of packets. To achieve low-latency video streaming, we propose a novel jitter management method named WMJitter using the watermark-based OOP mechanism. We mainly address the following challenges:
  • First, no existing studies have considered watermarks for jitter management in video conferencing systems. How can we use watermarks to handle disorder packets with low latency? (Section V-A)
  • Second, existing buffer-based jitter management adapt buffer size to trade-off buffering delay and packet loss. How can we achieve the same goal using watermarks? (Section V-B)
@@ V-B Balancing Latency with Packet Loss
The $UpdateWM()$ function uses an adaptive watermark generation algorithm to indicate the progress of data streams affected by network jitter. Upon the arrival of a new packet, we calculate an alternative watermark by subtracting the timestamp with a designated lag, subsequently selecting the larger value between the alternative watermark and the current watermark as the updated watermark. Formally, we have:
@@
Algorithm 2 CalculateLag($p$).
1:$p$: new arrived packet;
$jtHistory$: store network jitter for packets in a time window;
$histogram$: maintain the probability distribution of accumulated jitter in $jtHistory$; $PktSize$: the size of each packet.
2:$lag$
3:Calculate the transmitting jitter of packet $p$ compared to the last arrived packet;
4:Append $p$ to $jtHistory$ and update the window;
5:Update the $histogram$ of jitter according to $jtHistory$;
6:$lastTS=p.ts$;
7:if  $p.ts>latestTS$ then
8:    $disorder=0$ and $latestTS=p.ts$;
9:if  ! $disorder$ then
10:    $lag=histogram.Quantile(percentile))$;
11:else
12:    $i=argmin$ $Cost$, where $Cost=max(0,i-lag)+PktSize*100*(1-\sum_{k=0}^{i/PktSize}{histogram(k)})$;
13:    $lag=max(lag,i);$
14:return $lag$;
The $lag$ is a key parameter that directly impacts the balance between buffering delay and loss rate. A larger lag leads to increased buffering delay and reduced loss rate, while a smaller lag leads to an opposite scenario. Intuitively, a lag value that closely approximates the real-time jitter of transmitting latency optimizes the balance between buffer delay and loss rate. Thus, it is necessary to accurately evaluate the transmitting latency jitter to enable adaptive and timely watermark adjustment in Equation 3. To ensure precise evaluation of the jitter of transmitting latency, we employ a statistical method as shown in Algorithm 2.
Our approach utilizes a sliding window (typically set at 2 seconds) to retain the most recent arrival packets with their timestamps, enabling the calculation of jitter between the packets within the window. This is because packets with timestamps close to the current data packet are the most important for accurately estimating the network jitter. The algorithm employs two arrays $jtHistory$ and $histogram$, where $jtHistory$ maintains the jitter values of all data packets within a timestamp window and $histogram$ maintains the probability distribution of all accumulated jitter values obtained in history in the form of a histogram. The algorithm first determines if the current packet is out-of-order, and in the event of an in-order packet, it calculates the estimated network jitter as the value corresponding to a certain percentile (e.g., the 95th percentile) of the $histogram$ array and assigns it to $lag$ (Lines 7-8). Conversely, for out-of-order packets, the estimated network jitter is derived by identifying the value that minimizes a designated cost function (Lines 10-11).
The cost function (Line 10) considers both buffer delay and packet loss rate. The first term of the function indicates the influence of the estimated jitter on the packet’s delay in the buffer: the smaller the estimated network jitter, the lower the buffer delay. The second operator represents the effect of estimated jitter on the cost of packet loss: as the estimated network jitter approaches higher percentiles in the histogram, the probability of packet loss decreases and the expected cost of packet loss decreases. When the jitter value minimizes the cost function, it means that the lowest buffer delay is achieved while ensuring no packet loss.
This method ensures an accurate and real-time jitter assessment, which, in turn, informs the watermark updates to adapt to the dynamic network performance effectively.
@@ VI-A Experimental Setup
TABLE II: Location of endpoints and users
Provider
Endpoint Location
User Location
Tencent cloud
Tokyo, Guangzhou, Singapore, Frankfurt, Sao Paulo
Shenzhen, Singapore, Hong Kong
Wonder Proxy
Brazil, Australia, Singapore, South Africa, Greece
Argentina, Indonesia, Uganda, United States, Lithuania
Geo-distributed setting. We perform our experiments based on the network performance traces collected from Tencent cloud [12] and WonderProxy [24], as shown in Table II.
@@
Compared approaches. To show the superiority of our methods, we adopt three routing methods and two jitter management approaches for comparison. End-to-end latency is the main evaluation metrics for comparing different approaches.
@@
Configurations. For each geo-distributed setting, we simulate five meeting sessions hosted on the five endpoint servers. When one endpoint is selected to host a meeting, the other four endpoints are adopted as relays. As we only consider relay routes with less than two hops, there are 17 possible paths for each $<$endpoint, user$>$ pair. For each meeting, 600,000 packets are sent between each $<$endpoint, user$>$ pair to simulate a meeting session of 100 minutes long. During the packet handling, some late packets are dropped and cause packet loss. We report the loss rate and the end-to-end latency of successful packets as the evaluation metrics for all compared algorithms.
@@
VI-B Network Characteristics
We first study the characteristics of the two geo-distributed environments. As shown in Figure 8, which illustrates the average network performance and variance, the two platforms have very different features. First, the network performance between WonderProxy servers are highly dynamic, probably because there are many users utilizing the proxy services at the same time. In contrary, network performance between Tencent cloud servers and servers to clients are more stable, making the performance of relay routes more predictable. Second, comparing the average network latency between endpoint-endpoint and endpoint-user pairs, we found that there are some routes in Tencent cloud with extremely low latency. For example, the latency between $<$GZ, UserSZ$>$ is only 6ms, which is 1/50 of the latency between $<$GZ, UserSG$>$. In contrary, the network latency between different $<$source, destination$>$ pairs are less heterogeneous. The highest latency (i.e., 877ms) is around ten times as big as the lowest latency (i.e., 71ms).
Since the two platforms differ a lot on the network features, we believe our trace-driven evaluation results are representative to show the effectiveness of our proposed algorithms.
@@ VI-B Network Characteristics
VI-C Trace-Driven Evaluation Results
VI-C1 End-to-end latency on WonderProxy
@@
First, across all the meeting sessions hosted in different data centers, VCR-WM obtains the best end-to-end latency for all users distributed globally, while DRT-BF obtains the worst results in most cases. Take the results of Greece endpoint as an example. Figures 8(c) shows that, VCR-WM guarantees all packets to arrive within 640ms for users located in Uganda, which is even lower than the shortest end-to-end latency results obtained by DRT-BF. Compared to VIA-BF, VCR-WM reduces the average end-to-end latency by 4%-13%. Comparing different users in the same meeting session, VCR-WM guarantees that 50% of the packets experience a lag less than 375ms, 730ms, 500ms, 502ms and 72ms for clients located in Argentina, Indonesia, Uganda, United States and Lithuania, respectively. That is, users closer to the endpoint are more likely to obtain better video conferencing experiences. For example, Lithuania, which is also in Europe as Greece, is geographically the closest to the endpoint and thus has the lowest average end-to-end latency among all users. What’s more, all compared methods work similarly for Lithuania, since the direct route has much lower transmitting latency compared to the other relay routes. As a result, Direct, Via and VCRoute all select the same direct path. This further increases the stability of packet latency and reduces jitter.
@@
Second, to study the effectiveness of WMJitter, we compare the results of DRT-BF with those of DRT-WM and the results of VIA-BF with those of VIA-WM. As shown in Figure 7(a), the network performance between endpoints and users IDN, UGA and US are much more dynamic than the other two users. As a result, we have observed much higher improvement of the watermark-based jitter management methods compared to the buffer-based method on reducing the end-to-end latency. For example, DRT-WM reduces the end-to-end latency by 8%-12%, 9%-16% and 10%-16% compared to DRT-BF for clients in IDN, UGA and US, respectively, across the three meeting sessions. Similarly, VIA-WM reduces the end-to-end latency by 8%-13%, 2%-9% and 5%-10% compared to VIA-BF for clients in IDN, UGA and US, respectively, across the three meeting sessions. The above results have shown the effectiveness of our watermark-based jitter management.
Finally, comparing VCR-WM with DRT-WM and VIA-WM, different routing methods did not lead to much difference on the end-to-end latency results. This is mainly because that the network performance has low heterogeneity in WonderProxy. Thus, when selecting different routes, the resulted transmitting latency does not differ much. This is especially true when users are UserAR (Argentina) and UserLT (Lithuania), which have relatively lower endpoint-user network performance variances as shown in Figure 7(a). This explains the similar performance of different algorithms in Figures 9-11 when users are AR and LT. VCR-WM performs much better than DRT-WM for the meeting session in Greece when the user is UGA (Uganda), as shown in Figure 8(c). This is because the latency of the direct route between $<$GR, UserUGA$>$ is high (708ms) while there is a much lower relay route via ZA server, namely $<$GR, ZA, UserUGA$>$, that can effectively reduce the average transmitting latency to 355ms. This shows that VCRoute is smart enough to find good relay routes. Similar observation can also be found in Figure 10(c).
VI-C2 End-to-end latency on Tencent Cloud
@@
The evaluation results on Tencent cloud has shown some similar observations as those obtained from WonderProxy. Below we will only discuss the differences between Tencent cloud and WonderProxy.
@@
Figure 12 and 13 show the CDFs of end-to-end latency results of users obtained by VCR-WM, DRT-WM and VIA-WM when meetings are hosted on the Tokyo and Sao Paulo servers. We deliberately removed the results of DRT-BF and VIA-BF from the figures since they performed very closely to DRT-WM and VIA-WM due to the low variances of the network performances. Different from the results on WonderProxy, VCR-WM obtains much lower end-to-end latency compared to VIA-WM and DRT-WM. For example, the 50%-percentile latency of VCR-WM is 0%-16% and 6%-18% lower than those of DRT-WM and VIA-WM, respectively. This demonstrates the effectiveness of VCRoute on finding better routes than other state-of-the-art routing schedulers for video conferencing systems.
@@
VI-C3 Loss rate results
We further study the loss rate of video streams resulted from different comparisons. Figure 15 shows the loss rate results per meeting session on both WonderProxy and Tencent cloud. As expected, the loss rate on WonderProxy is in general much larger than that on Tencent cloud, due to the much higher network performance variances on WonderProxy. The loss rates of meeting sessions using Direct routing on Tencent cloud are almost zero since there’s no significant jitter on the route. Thus no packet arrives too late to be dropped. In contrary, the loss rate is high even on direct routes on WonderProxy due to the high network performance variances. Further, watermark-based jitter management method can slightly reduce the loss rate compared to buffer-based methods. For example, DRT-WM and VIA-WM reduce the loss rate by 2%-3% and 2%-4% compared to DRT-BF and VIA-BF, respectively, on WonderProxy. The Via routing method (i.e., VIA-BF and VIA-WM) generated much higher loss rates on Tencent cloud than the other methods, due to its tendency to explore which leads to more path changes. For example, there are 34,020 times of path changes with Via for the $<$Tokyo, Shenzhen$>$ data stream with 600,000 packets. Although VCRoute also explores during the search for a good route, it deliberately reduces the number of path changes to reduce network jitter. In all cases, VCR-WM yields lower or comparable packet loss rates compared to the other methods, while providing much lower end-to-end latency.
TABLE III: RMSE of our DNN prediction model on WonderProxy. Values in brackets are mean end-to-end latency (ms).
UserAR
UserIDN
UserUGA
UserUS
UserLT
GR
23(383)
126(755)
57(514)
95(510)
76(123)
ZA
57(474)
135(796)
3(98)
159(738)
15(239)
SG
14(426)
134(481)
4(412)
84(661)
45(362)
TABLE IV: RMSE of our DNN prediction model on Tencent. Values in brackets are mean end-to-end latency (ms).
UserSZ
UserSG
TYO
1(137)
12(97)
SP
0.5(353)
10(362)
Figure 14: Relative error distribution of end-to-end latency prediction between $<$ZA,UserIDN$>$ DCs.
VI-C4 Accuracy of end-to-end latency prediction
To show how our DNN-based prediction model helps to obtain the low end-to-end latency results on both clouds, we present the RMSE values of predicted latency of packets between different pairs of DCs in Table III and IV. Values in brackets represent the mean values of actual end-to-end latency results. Overall, the prediction error on Tencent is much lower than that on WonderProxy. This is because the network performance is highly dynamic on WonderProxy. When inferring the end-to-end latency using our prediction model, the input data (i.e., transmitting latency and jitter) are the latest measurements from the system and can be largely different from the real network performance experienced by the current packet. Nevertheless, although the average error looks big for highly dynamic paths (e.g., RMSE=135 for $<$ZA,UserIDN$>$), our model generates good prediction in most cases. For example, Figure 14 shows the relative error distribution of end-to-end latency prediction for packets between ZA and UserIDN. Over 80% of the predictions have relative error below 10%.
@@ VI-C4 Accuracy of end-to-end latency prediction
VI-C5 System overhead
With our top-k pruning, the overhead of VCRoute searching for the best route is very low ( 0.3ms for each packet). Due to the centralized design of VCRoute, the main overhead introduced by our system to video streams comes from the extra time needed to transfer routing decisions from endpoints to users. Figure 16 shows the additional overhead per packet caused by transmitting routing decisions for each meeting session hosted on different endpoints in WonderProxy and Tencent. Similar to our observations above, the additional overhead is larger for users with higher network variances (e.g., UserIDN on WonderProxy and UserSG on Tencent). The largest overhead for UserIDN is 9.1ms, which is neglectable compared to the end-to-end latency per packet as shown in Figure 9(b). The results show that, our design introduces an additional overhead less than 1% of the average end-to-end latency for each packet.
@@ VI-C5 System overhead
Figure 16: Additional overhead per packet caused by VCRoute
VI-D Scalability Study
To study the effectiveness of our system under larger geo-distributed environments, we constructed three synthetic network environments with one user communicating with one endpoint, and there are 9, 90 and 900 relay paths between them. For each path, we set the average latency to a random number between [100ms, 200ms]. One third of the paths have standard deviation of 10, one third have 20 and the rest have 30. Compared to the state-of-the-art VIA-BF, VCR-WM reduces the average end-to-end latency by 27%, 29% and 29% at scale 9, 90 and 900, respectively. This demonstrates that VCR-WM can achieve good optimization effectiveness even under a larger system scale. Figure 17 shows the CDFs of the end-to-end latency results.
(a) #paths=9
(b) #paths=90
(c) #paths=900
Figure 17: End-to-end latency under different system scales.
VI-E Discussions
@@
  • VCRoute is effective for networks with highly heterogeneous networks, as demonstrated by the results on Tencent cloud.
@@ VIII Conclusion
In this paper, we have addressed two significant challenges in optimizing video conferencing systems across geo-distributed data centers. First, we identified the limitations of existing routing methods that primarily focus on minimizing packet transmitting latency, which doesn’t necessarily translate to low end-to-end latency. In response, we introduced VCRoute, an application-specific packet routing technique. VCRoute incorporates both network transmission time and packet reordering time, effectively reducing the end-to-end latency. Second, we observed that the inter-DC network latency can drastically fluctuate, resulting in high jitter detrimental to video conferencing applications. Traditional buffer-based jitter management methods often introduce unnecessary delays, especially when handling stragglers. To overcome this, we proposed WMJitter, a watermark-based Out-Of-Order processing mechanism tailored to manage jitter at the users’ end. Our experimental evaluation, conducted in two distinct real-world geo-distributed environments, demonstrates the efficacy and viability of our proposed solutions, VCRoute and WMJitter, for enhancing the performance of video conferencing systems in such complex settings.

### v2

@@ V Watermark-based Jitter Manager
The jitter manager sits on packet receivers’ side to ensure ordered emit of packets. To achieve low-latency audio streaming, we propose a novel jitter management method named WMJitter using the watermark-based OOP mechanism. We mainly address the following challenges:
  • No existing studies have considered watermarks for jitter management in video conferencing systems. How can we use watermarks to handle disorder packets with low latency? (Section V-A)
  • Existing buffer-based jitter management adapt buffer sizes to trade-off buffering delay and packet loss. How can we achieve the same goal using watermarks? (Section V-B)
@@ V-B Balancing Latency with Packet Loss
The $UpdateWM()$ function uses an adaptive watermark generation algorithm to indicate the progress of data streams affected by network jitter. Upon the arrival of a new packet, we calculate an alternative watermark by subtracting the timestamp with a designated lag, subsequently selecting the larger value between the alternative watermark and the current watermark as the updated watermark. Formally:
@@
The $lag$ is a key parameter that directly impacts the balance between buffering delay and loss rate. A larger lag leads to increased buffering delay and reduced loss rate, while a smaller lag leads to an opposite scenario. Intuitively, a lag value that closely approximates the real-time jitter of transmitting latency optimizes the balance between buffer delay and loss rate. To enable adaptive and timely watermark adjustment, we employ the window-based statistical method of estimating jitter used in WebRTC NetEQ module [12] to determine the $lag$ value.
@@ VI-A Experimental Setup
TABLE I: Location of endpoints and users
Provider Endpoint Location User Location
Tencent cloud Tokyo, Guangzhou, Singapore, Frankfurt, Sao Paulo Shenzhen, Singapore, Hong Kong
Wonder Proxy Brazil, Australia, Singapore, South Africa, Greece Argentina, Indonesia, Uganda, United States, Lithuania
Geo-distributed setting. We perform our experiments based on the network performance traces collected from Tencent cloud [16] and WonderProxy [27], as shown in Table I.
@@
As shown in Figure 7, which illustrates the average network performance and variance, the two platforms have very different features. First, the latency between WonderProxy servers are higher and more dynamic than Tencent cloud. Second, the network latency between different $<$source, destination$>$ pairs of WonderProxy are less heterogeneous. Since the two platforms differ a lot on the network features, we believe our trace-driven evaluation results are representative to show the effectiveness of our proposed algorithms.
Compared approaches. To show the superiority of our methods, we adopt three routing methods and two jitter management approaches for comparison. End-to-end latency is the main metric for comparing different approaches.
@@
Configurations. For each geo-distributed setting, we simulate five meeting sessions hosted on the five endpoint servers. When one endpoint is selected to host a meeting, the other four endpoints are adopted as relays. As we only consider relay routes with less than two hops, there are 17 possible paths for each $<$endpoint, user$>$ pair. For each meeting, 600,000 packets are sent between each $<$endpoint, user$>$ pair to simulate a meeting session of 100 minutes long. During the packet handling, some late packets are dropped and cause packet loss. We report the loss rate and the end-to-end latency of successful packets as evaluation metrics.
@@
VI-B Trace-Driven Evaluation Results
VI-B1 End-to-end latency on WonderProxy
@@
First, across all the meeting sessions hosted in different data centers, VCR-WM obtains the best end-to-end latency for all users distributed globally, while DRT-BF obtains the worst results in most cases. Take the results of Greece endpoint as an example. Figures 8(c) shows that, VCR-WM guarantees all packets to arrive within 640ms for users located in Uganda, which is even lower than the shortest end-to-end latency results obtained by DRT-BF. Compared to VIA-BF, VCR-WM reduces the average end-to-end latency by 4%-13%. Note that, for some cases, the VCR-WM and DRT-WM obtains similar end-to-end latency results (i.e., $<$ZA, UserAR$>$). This is because that the best relay path is the direct path, and the transmitting latency is far less than the other paths. Similarly, for some cases, the VCR-WM and VIA-WM obtains similar end-to-end latency results (i.e., $<$SG, UserUGA$>$) because the best path is far better than others. Comparing different users in the same meeting session, users closer to the endpoint are more likely to obtain better video conferencing experiences. For example, Lithuania, which is also in Europe as Greece, is geographically the closest to the endpoint and thus has the lowest average end-to-end latency among all users. What’s more, all compared methods work similarly for Lithuania, since the direct route has much lower transmitting latency compared to the other relay routes. As a result, Direct, Via and VCRoute all select the same direct path. This further increases the stability of latency and reduces jitter.
@@
Second, to study the effectiveness of WMJitter, we compare the results of DRT-BF with those of DRT-WM and the results of VIA-BF with those of VIA-WM. As shown in Figure 7(a), the network performance between endpoints and users IDN, UGA and US are much more dynamic than the other two users. As a result, we have observed much higher improvement of the watermark-based jitter management methods compared to the buffer-based method on reducing the end-to-end latency. For example, DRT-WM reduces the end-to-end latency by 8%-12%, 9%-16% and 10%-16% compared to DRT-BF for clients in IDN, UGA and US, respectively, across the three meeting sessions. Similarly, VIA-WM reduces the end-to-end latency by 8%-13%, 2%-9% and 5%-10% compared to VIA-BF for clients in IDN, UGA and US, respectively, across the three meeting sessions. The above results have demonstrated the effectiveness of WMJitter.
VI-B2 End-to-end latency on Tencent Cloud
@@
Evaluation results on Tencent cloud has shown similar observations as those from WonderProxy. Below we only discuss the differences between Tencent cloud and WonderProxy.
@@
Figure 11 and 12 show the CDFs of end-to-end latency results of users obtained by VCR-WM, DRT-WM and VIA-WM when meetings are hosted on the Tokyo and Sao Paulo servers. We deliberately removed the results of DRT-BF and VIA-BF from the figures since they performed very closely to DRT-WM and VIA-WM due to the low variances of the network performances. On these two cases, VCR-WM obtains much lower end-to-end latency compared to VIA-WM and DRT-WM. For example, the 50%-percentile latency of VCR-WM is 0%-16% and 6%-18% lower than those of DRT-WM and VIA-WM, respectively. This demonstrates the effectiveness of VCRoute on finding better routes than other state-of-the-art routing schedulers for video conferencing systems.
VI-B3 Loss rate results
We further study the loss rate of audio streams resulted from different comparisons. Figure 13 shows the loss rate results per meeting session on both WonderProxy and Tencent cloud. As expected, the loss rate on WonderProxy is in general much larger than that on Tencent cloud, due to the much higher network performance variances on WonderProxy. The loss rates of meeting sessions using Direct routing on Tencent cloud are almost zero since there’s no significant jitter on the route. Thus no packet arrives too late to be dropped. In contrary, the loss rate is high even on direct routes on WonderProxy due to the high network performance variances. Further, watermark-based jitter management method can slightly reduce the loss rate compared to buffer-based methods. For example, DRT-WM and VIA-WM reduce the loss rate by 2%-3% and 2%-4% compared to DRT-BF and VIA-BF, respectively, on WonderProxy. The Via routing method (i.e., VIA-BF and VIA-WM) generated much higher loss rates on Tencent cloud than the other methods, due to its tendency to explore which leads to more path changes. For example, there are 34,020 times of path changes with Via for the $<$Tokyo, Shenzhen$>$ data stream with 600,000 packets. Although VCRoute also explores during the search for a good route, it deliberately reduces the number of path changes to reduce network jitter. In all cases, VCR-WM yields lower or comparable packet loss rates compared to the other methods, while providing much lower end-to-end latency.
@@ VI-B3 Loss rate results
VI-B4 System overhead
With our top-k pruning, the overhead of VCRoute searching for the best route is very low ($\sim$0.3ms for each packet). Due to the centralized design of VCRoute, the main overhead introduced by our system to audio streams comes from the extra time needed to transfer routing decisions from endpoints to users. Figure 14 shows the additional overhead per packet caused by transmitting routing decisions for each meeting session hosted on different endpoints in WonderProxy and Tencent. Similar to our observations above, the additional overhead is larger for users with higher network variances (e.g., UserIDN on WonderProxy and UserSG on Tencent). The largest overhead for UserIDN is 9.1ms, which is neglectable compared to the end-to-end latency per packet as shown in Figure 9(b). It means that our design introduces an additional overhead less than 1% of the average end-to-end latency for each packet.
@@ VI-B4 System overhead
Figure 14: Additional overhead per packet of VCRoute
VI-C Discussions
@@
  • VCRoute is effective for DCs with highly heterogeneous network, as demonstrated by the results on Tencent cloud.
@@ VIII Conclusion
In this paper, we address two challenges for video conferencing systems across geo-distributed DCs. First, existing routing methods mainly focus on minimizing packet transmitting latency, which does not necessarily translate to low end-to-end latency. In response, we introduced VCRoute, an application-specific packet routing technique that jointly considers network transmission time and packet reordering time to reduce end-to-end latency. Second, inter-DC network latency can drastically fluctuate, resulting in high jitter detrimental to video conferencing applications. Traditional buffer-based jitter management methods often introduce unnecessary delays, especially when handling stragglers. To overcome this, we proposed WMJitter, a watermark-based Out-of-Order Processing mechanism tailored to manage jitter at the users’ end. Evaluations using two distinct real-world geo-distributed environments have demonstrated the efficacy and viability of our proposed solutions on enhancing the performance of video conferencing systems.